The product of a nonzero integer and –1 has same value as the square of the integer. What is the integer?
(A) –2 (B) –1 (C) 0 (D) 1 (E) 2

Answer by richwmiller(17219)

Since we weren't told how we had to solve this (trial and error or algebraically)
we can try -2 then -1, 0, 1 and 2
A)
-1*-2=2
2^2=4
2 doesn't equal 4
A is not the answer
nor is E)
-1*2=-2
-2*-2=4
B)
-1*-1=1
-1*-1=1
they are the same
Answer is B) -1
C)
-1*0=0
0*0=0
They are the same
C) is the answer
No it said non zero integer!!!
C) is not the answer
D)
-1*1=-1
1^2=1
they are not the same
E)
-1*2=-2
2^2=4
they are not the same
Now the algebraic solution
-1*x=x^2
divide by x
-1=x
the non zero solution is -1
Something to remember is that x can be zero too but not in this problem.